OpenStudy (anonymous):

Solve the quadratic equation 5x2 - 125 = 0. A. -5 B. 5 C. ±5 D. ±25

OpenStudy (stamp):

hint: factor your expression

OpenStudy (ryan123345):

C because when you plug in the numbers it equals 0 :)

OpenStudy (stamp):

lol but what if the multiple choice wasn't there...

OpenStudy (stamp):

is it game over? because guess and check isnt a good way to "learn" math

OpenStudy (ryan123345):

you arent funny

OpenStudy (stamp):

sounds like a headache, why not just factor it and set terms equal to zero

OpenStudy (e.cociuba):

Wouldn't there be 2 answers here 5 and -5?

OpenStudy (stamp):

\[5x^2-125=0\]\[5(x^2-25)=0\]\[5(x+5)(x-5)=0\] Identify and solve (x + 5) = 0 (x - 5) = 0 x = +/- 5 ps you are not funny either

OpenStudy (e.cociuba):

Its C.)

OpenStudy (stamp):

@e.cociuba The answer choice includes a +/- sign, so +/- 5 = +5, -5

OpenStudy (e.cociuba):

Opps srry lol I didn't see tht.